SME Company has a debt-equity ratio of .70. Return on assets Is 7 percent, and total equity is $481,000. a. What is the equity multiplier? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) What is the return on equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) What is the net income? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.) b. c. a. Equity multiplier b. Return on equity c. | Net income times
Answer #1

Debt to Equity Ratio = 0.7

Return on Assets = 7%

Equity = $481,000

So,

Total Debt = Equity * debt to equity ratio

Total Debt = 481000 * 0.7

Debt = $336,700

Total Assets = Debt + Equity

Total Assets = 336700 + 481000

Total Assets = $817,700

Net Profit = ROA * Total Assets

Net Profit = 0.07 * 817700

Net Profit = $57,239

Equity Multiplier = Total Assets/Total Equity

Equity Multiplier = 817700/481000

Equity Multiplier = 1.7

Return on Equity = Net Profit/Equity

Return on Equity = 57239/481000

Return on Equity = 11.9%

Net Income = $57,239